OpenStudy (anonymous):

Solve by elimination 9x + 2y = 2 21x + 6y = 4

OpenStudy (anonymous):

can i multiply the first equation by a -3?

OpenStudy (anonymous):

yeah

OpenStudy (aric200):

x= 1 over 3 y= -1 over 2

OpenStudy (anonymous):

-27x-6y=-6 21x +6y= 4 ----------- -6x =-2 x = 1/3

OpenStudy (aric200):

sorry its x= -1 half

OpenStudy (anonymous):

Good job @Reaper534

OpenStudy (aric200):

negative 1 half

OpenStudy (anonymous):

ahh ok i see then i plug in x to the 2nd equation haha thanks

OpenStudy (anonymous):

Thanks guys